OpenStudy (anonymous):

What is the missing exponent? (6 )^-4 = 6^20

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@skullpatrol

OpenStudy (skullpatrol):

$$\Huge (a^m)^n=a^{mn}$$

OpenStudy (anonymous):

So 20?

OpenStudy (anonymous):

-20?

OpenStudy (skullpatrol):

$$\Huge (6^m)^{-4}=6^{20}$$

OpenStudy (anonymous):

So 20?

OpenStudy (skullpatrol):

$$\Huge m \cdot {(-4)} = 20$$

OpenStudy (anonymous):

Oh my derp. 24.

OpenStudy (skullpatrol):

a raised dot means multiply

OpenStudy (anonymous):

5?

OpenStudy (skullpatrol):

what about the sign?

OpenStudy (anonymous):

-5

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

ok i got 1 more